The measure of the supplement of an angle is eighty more than the measure of the angle. Find the measure of both angles.
Answer by edjones(8007)

Let s=supplement and a=angle
s=a+80 for this angle
s=180-a for any angle
a+80=180-a (s=s)
2a=260
a=130 deg
s=50 deg
.
Ed